OpenStudy (anonymous):

Find an equation for the parabola with focus at (10, -4) and vertex at (3,-4) Answer = y^2 + 8y - 28x + 100 = 0

OpenStudy (anonymous):

the distance between the vertex and the focus is the quantity a in 4a(y-k)=(x-h)^2 where (h,k) is the vertex. h=3 k=-4 so just find the distance between the vertex and focus. and plug in all those values.

OpenStudy (anonymous):

ooh.. after that put in general form.. :)

OpenStudy (campbell_st):

the line of symmetry is y = -4 so the parabola is concave right and has the general form is \[4a(x -h) = (y - k)^2 \] (h, k) is is the vertex and a = focal length which gives a = 7 and since the focus is above the verex the parabola is concave up \[4 \times 7(x -3) = (y + 4)^2\] expand and simplify will give your answer

OpenStudy (anonymous):

And there are 3 other forms, correct?